# Mitigating the Backdoor Effect for Multi-Task Model Merging via Safety-Aware Subspace

### v1

@@ 4.2 Experimental Results
DAM can achieve comparable or better effects in addressing the backdoor issues without additional training compared with post-defense methods. Through the comparison experiments among Concrete AM(layer-wise), post-defense methods(ANP, AWM, and SAU), and DAM shown in Table 2 and Table 9, we observe that though previous post-defense methods can further mitigate the backdoor issues on the SOTA merged model in a way, they are still clearly worse than DAM in addressing the backdoor issues. The poor effectiveness of these post-defense methods can be attributed to their reliance on high-quality labeled data (Wu et al., 2024), which may be unrealistic during merging. Additionally, their target is a merged model that has converged solely on performance, making it difficult to achieve a balance between performance and safety in the post-defense process. Notably, these post-defense methods introduce additional training costs but DAM can naturally cope with the backdoor issues in an end-to-end training manner without this constraint.
@@
[…]
DAM can achieve better results than other backdoor-related merging methods that are not designed for multi-task merging scenarios or successfully defend their proposed backdoor attack for model merging. Both WAG (Arora et al., 2024) and LoRA-as-an-Attack (Liu et al., 2024) defend the backdoor by directly averaging the homogeneous clean and backdoored full model weights or LoRa on the same task without other complex designs(e.g.subspace). To distinguish it from our proposed DAM for multi-task merging, as the experiments show in Table 6, for each task related to the backdoored individual finetuned model, we additionally select a clean model for this task during multi-task merging. This corresponds to the open-source community scenario, where exists many models for the same task, and part of them are injected with backdoors but we can not know in advance. From the results, we can observe that DAM consistently achieves higher ACC and lower ASR in different settings. For the BadMering (Arora et al., 2024), which aims to introduce a new backdoor attack to break the safeguard for existing merging methods. We mainly utilize it to attack the merged model of Table 2 and Table 3 adopting the DAM strategy. The results of Table 6 show that it’s difficult to clearly increase the ASR adopting the attack proposed by BadMerging. In other words, DAM can successfully defend this new backdoor attack, further verifying its effectiveness in addressing the backdoor issues.
Table 5: Comparison results with other backdoor-defense methods that are not designed for multi-task scenarios.
@@
Two masks have different effects on DAM and collectively promote the merged model with a good trade-off between safety and performance. The experimental settings are consistent with Table 2 and Table 3. As shown in Eq.3, removing the $M_{2}$ of DAM means we only focus on the interference of multi-task merging, which is similar to Concrete (Tang et al., 2023). In contrast, removing the $M_{2}$ of DAM means we mainly deal with the backdoor issues, which can be achieved by adopting AdaMerging (Yang et al., 2023) on the perturbated data. Then, removing two masks together means we only focus on learning the merging coefficients without revising the task vectors. The results shown in Table 7 can verify the effectiveness of the dual-mask optimization in addressing the backdoor issues for multi-task merging.
@@ A.1 Multi-Task Model Merging
Multitask model merging aims to provide cost-effective parameter fusion strategies to integrate multiple task-specific finetuned models from the shared pre-trained model into a unified one that can handle various tasks (Tang et al., 2024; Yang et al., 2024a). The well-known strategy to merge multiple task-specific models is by performing element-wise interpolation on weights, such as Weight Average (Wortsman et al., 2022; Kaddour, 2022; Chronopoulou et al., 2023; Sanyal et al., 2023; Lawson & Qureshi, 2024; Jia et al., 2024), Fisher merging (Matena & Raffel, 2022; Ryu et al., 2023; Nathan et al., 2024), RegMean (Jin et al., 2022), and Task Arithmetic (Ilharco et al., 2022a; Zhang et al., 2023; Ni et al., 2023; Ortiz-Jimenez et al., 2024). To further enhance the effectiveness of merging, many efforts have been devoted to addressing interference among tasks and proposing gradient-conflict-based (Yadav et al., 2024), representation-based (Yang et al., 2023; 2024b), and subspace-based methods (Tang et al., 2023; Tam et al., 2024; Yu et al., 2024). However, these works aim to facilitate knowledge transfer during model merging, but harmful elements like adversarially injected backdoors may also succeed and transfer. Unfortunately, there is insufficient analysis of how multi-task merging will be affected if one or more task-specific models are injected with the backdoor.
@@ A.2 Model Merging Considering the Safety
The safety of machine learning algorithms is the key to their widespread applications (Wu et al., 2024). Recently, many researchers have begun to emphasize safety concerns associated with merging scenarios (Yang et al., 2024a). For example, some works adopt the subspace-based and data-aware merging methods to deal with the misalignment (Yi et al., 2024; Hammoud et al., 2024) and Intellectual Property Protection (Cong et al., 2024) problems for large language models (LLMs). Consistent with our work that focuses on the backdoor issues, WAG (Arora et al., 2024) and LoRA-as-an-Attack (Liu et al., 2024) conducted preliminary exploration under backdoor defense scenarios. However, they only adopt the initial weight average strategy (Wortsman et al., 2022), to cope with the backdoor issues on the same task without more fine-grained designs(e.g. subspace). Besides, BadMerging (Zhang et al., 2024) only focuses on executing effective backdoor attacks to maximize the ASR to break the safeguard for multi-task merging methods, while we mainly consider the defense strategy to minimize the ASR during merging. To the best of our knowledge, our work is the first to conduct extensive experiments to investigate the backdoor effect for multi-task merging scenarios and provide a novel defense-aware merging algorithm to alleviate this problem.
@@ B.2 The Pseudo Code of Our Proposed DAM
Corollary 1 (Performance-safety trade-off control).
@@
6:     Mask task vectors $\mathcal{T}$ with $\mathbf{m}$ to get $\mathcal{T}^{\prime}$
7:     Rescale the masked task vectors $\mathcal{T}^{\prime}$ to get $\mathcal{T}^{\prime\prime}$
@@
8:     Initialize merging coefficient $\lambda=\{\lambda_{i}\}_{i=1}^{n}$ associated with model merging algorithm
9:     $\theta\leftarrow\text{MergeWeight}(\theta_{\text{pre}},\mathcal{T}^{\prime\prime};\lambda)$
10:     if $\lambda$ is optimizable then
11:        for each task $s_{i}\in\mathcal{S}^{\text{test}}$ do
12:           Sample a batch of unlabeled data $\mathcal{D}_{i}$ from $s_{i}$
13:           $l_{i}\leftarrow\mathcal{L}_{i}(f(\theta),\mathcal{D}_{i})$
14:           $l_{i}^{perturbation}\leftarrow\mathcal{L}_{i}(f(\theta),\mathcal{D}_{i}+\Delta_{i})$
15:           Clip $\Delta_{i}$: $\Delta_{i}=\Delta_{i}\times\min(1,\frac{\xi}{\|\Delta_{i}\|_{1}})$
16:        end for
17:        $\lambda^{\prime}\leftarrow\lambda-\eta_{1}\nabla_{\lambda}\left(\sum_{i=1}^{n}l_{i}\right)$
@@
18:        $\theta\leftarrow\text{MergeWeight}(\theta_{\text{pre}},\mathcal{T}^{\prime\prime};\lambda^{\prime})$ // Update the merged model with the updated $\lambda^{\prime}$
@@
19:        $\Delta\leftarrow\Delta-\eta_{2}\nabla_{\Delta}\left(\sum_{i=1}^{n}l_{i}^{perturbation}\right)$ // Adversarial Trigger Recovery
20:     end if
21:     for each task $s_{i}\in\mathcal{S}$ do
22:        Sample a batch of unlabeled data $\mathcal{D}_{i}$ from $s_{i}$
23:        $l_{i}\leftarrow\mathcal{L}_{i}(f(\theta),\mathcal{D}_{i})$
24:        $l_{i}^{perturbation}\leftarrow\mathcal{L}_{i}(f(\theta),\mathcal{D}_{i}+\Delta_{i})$
25:     end for
26:     $\mathbf{x}\leftarrow\mathbf{x}-\eta_{3}\nabla_{\mathbf{x}}\left(\sum_{i=1}^{n}(l_{i}+\alpha l_{i}^{perturbation})\right)$
27:  end for
28:  Return: the mask parameterized by logits $\mathbf{x}$.
@@ C.2 More Experiments about Backdoor Succession and Backdoor Transfer
From the results, we can observe that the evaluation of existing merging methods is not fully consistent with the results shown in (Tang et al., 2024) when we take performance and safety into consideration simultaneously. For example, Ties-Merging achieved unexpectedly poor results, sometimes even worse than RegMean and Fisher Merging, which is the opposite of its reported outcome. This can be attributed to the fact that the success of Ties-Merging relies heavily on the gradient conflict analysis of different tasks. However, this analysis just considers the task-specific performance without dealing with safety issues. When there exists backdoored task-specific models during merging, the causes of gradient conflicts are complex and multidimensional. Notably, compared with task-wise methods, Adamerging(task-wise) and Concrete AM(task-wise), the layer-wise version of corresponding methods can consistently achieve better performance, but their ASR results are still high. This further clarifies the single task-wise performance perspective is not always appropriate, highlighting the need for more exploration of the backdoor issues during multi-task merging.
@@ C.3 More Merging Results
Moreover, Table 10 and Table 11 display the merging results while merging two backdoored models and four clean models. Different from the Table 2 in the main text, we select other task-specific backdoored models during merging. These results can further verify the effectiveness and robustness of the proposed DAM.

### v2

@@ 4.2 Experimental Results
DAM can achieve comparable or better effects in addressing the backdoor issues without additional training compared with post-defense methods. Through the comparison experiments among Concrete AM (layer-wise), post-defense methods (ANP, AWM, and SAU), and DAM shown in Table 2 and Table 10, we observe that though previous post-defense methods can mitigate the backdoor issues on the SOTA merged model in a way, they are still clearly worse than DAM. This can be attributed to their reliance on high-quality labeled data (Wu et al., 2022), which is usually unrealistic during merging. Additionally, their operated target is a merged model that has converged solely based on performance, making it difficult to achieve a balance between performance and safety. Notably, as the efficiency studies of Table 8 show in the Appendix C.4, these post-defense methods introduce additional training costs but DAM can naturally cope with the backdoor issues in an end-to-end training manner without this constraint.
@@
DAM can achieve better results than other backdoor-related merging methods that are not designed for multi-task merging scenarios or successfully defend their proposed backdoor attack for model merging. To distinguish WAG (Arora et al., 2024) and LoRA-as-an-Attack (Liu et al., 2024) from our proposed DAM, as shown in Table 5, for each task related to the backdoored individual finetuned model, we additionally select a clean model for this task during multi-task merging. This corresponds to the open-source community scenario, where exists many models for the same task, and part of them are injected with backdoors but we can not know in advance. We can observe that DAM consistently achieves higher ACC and lower ASR in different settings. For the backdoor attack called BadMering (Arora et al., 2024), we mainly utilize it to attack the merged model of Table 2 and Table 3 adopting the DAM strategy. Combining DAM with BadMerging means we would like to check whether BadMerging can further inject backdoor-related parameters that DAM can not identify, further increasing the ASR badly. The results of Table 5 show that it’s difficult to clearly increase the ASR adopting the attack proposed by BadMerging. In other words, DAM can successfully defend this new backdoor attack, further verifying its effectiveness in addressing the backdoor issues.
Table 4: Comparison with other backdoor-defense methods that are not designed for multi-task scenarios.
@@
Two masks have different effects on DAM and collectively promote the merged model with a good trade-off between safety and performance. The experimental settings are consistent with Table 2 and Table 3. As shown in Eq.3, removing the $M_{2}$ of DAM means we only focus on the interference of multi-task merging, which is similar to Concrete (Tang et al., 2023). In contrast, removing the $M_{2}$ of DAM means we mainly deal with the backdoor issues, which can be achieved by adopting AdaMerging (Yang et al., 2023) on the perturbated data. Then, removing two masks together means we only focus on learning the merging coefficients without revising the task vectors. The results shown in Table 6 can collectively verify the effectiveness of the dual-mask optimization of DAM.
@@ 5 Conclusion
ACKNOWLEDGMENTS
We sincerely thank the anonymous reviewers for their valuable feedback, which has helped us polish the paper. This work is supported by STI 2030—Major Projects (No. 2021ZD0201405), Shenzhen Basic Research Project (Natural Science Foundation) Basic Research Key Project (NO. JCYJ20241202124430041), the Major Project in Judicial Research from Supreme People’s Court (NO. GFZDKT2024C08-3), and Huawei AI 100 Schools Program.
@@ A.1 Multi-Task Model Merging
Multitask model merging aims to provide cost-effective parameter fusion strategies to integrate multiple task-specific finetuned models from the shared pre-trained model into a unified one that can handle various tasks (Tang et al., 2024a; Yang et al., 2024a). The well-known strategy to merge multiple task-specific models is by performing element-wise interpolation on weights, such as Weight Average (Wortsman et al., 2022; Kaddour, 2022; Chronopoulou et al., 2023; Sanyal et al., 2023; Lawson & Qureshi, 2024; Jia et al., 2024), Fisher merging (Matena & Raffel, 2022; Ryu et al., 2023; Nathan et al., 2024), RegMean (Jin et al., 2022), and Task Arithmetic (Ilharco et al., 2022a; Zhang et al., 2023; Ni et al., 2023; Ortiz-Jimenez et al., 2024). To further enhance the effectiveness of merging, many efforts have been devoted to addressing interference among tasks and proposing gradient-conflict-based (Yadav et al., 2024), representation-based (Yang et al., 2023; 2024b), routing-based (Zhao et al., 2024a; b; Tang et al., 2024c; b), and subspace-based methods (Tang et al., 2023; Tam et al., 2024; Yu et al., 2024). Unfortunately, existing model merging studies enhance knowledge transfer but neglect adversarial backdoor propagation through parameter fusion, with insufficient analysis of multi-task system vulnerabilities.
@@ A.2 Model Merging Considering the Safety
The safety of machine learning algorithms is the key to their widespread applications (Wu et al., 2022). Recently, many researchers have begun to emphasize safety concerns associated with merging scenarios (Yang et al., 2024a). For example, some works adopt the subspace-based and data-aware merging methods to deal with the misalignment (Yi et al., 2024; Hammoud et al., 2024) and Intellectual Property Protection (Cong et al., 2024) problems for large language models (LLMs). Consistent with our work that focuses on the backdoor issues, WAG (Arora et al., 2024) and LoRA-as-an-Attack (Liu et al., 2024) conducted preliminary exploration under backdoor defense scenarios. However, they only adopt the initial weight average strategy (Wortsman et al., 2022), to cope with the backdoor issues on the same task without more fine-grained designs(e.g. subspace). Besides, BadMerging (Zhang et al., 2024) only focuses on executing effective backdoor attacks to maximize the ASR to break the safeguard for multi-task merging methods, while we mainly consider the defense strategy to minimize the ASR during merging. To the best of our knowledge, our work is the first to conduct extensive experiments to investigate the backdoor effect for multi-task merging scenarios and provide a novel defense-aware merging algorithm to alleviate this problem.
@@ B.2 The Pseudo Code of Our Proposed DAM
Proof of Theorem 1.
@@
6:     Mask task vectors $\mathcal{T}$ with $\mathbf{m}$ to get $\mathcal{T}^{\prime}$, Rescale the masked task vectors $\mathcal{T}^{\prime}$ to get $\mathcal{T}^{\prime\prime}$
@@
7:     Initialize merging coefficient $\lambda=\{\lambda_{i}\}_{i=1}^{n}$ associated with model merging algorithm
8:     $\theta\leftarrow\text{MergeWeight}(\theta_{\text{pre}},\mathcal{T}^{\prime\prime};\lambda)$
9:     if $\lambda$ is optimizable then
10:        for each task $s_{i}\in\mathcal{S}^{\text{test}}$ do
11:           Sample a batch of unlabeled data $\mathcal{D}_{i}$ from $s_{i}$
12:           $l_{i}\leftarrow\mathcal{L}_{i}(f(\theta),\mathcal{D}_{i})$
13:           $l_{i}^{perturbation}\leftarrow\mathcal{L}_{i}(f(\theta),\mathcal{D}_{i}+\Delta_{i})$
14:           Clip $\Delta_{i}$: $\Delta_{i}=\Delta_{i}\times\min(1,\frac{\xi}{\|\Delta_{i}\|_{1}})$
15:        end for
16:        $\lambda^{\prime}\leftarrow\lambda-\eta_{1}\nabla_{\lambda}\left(\sum_{i=1}^{n}l_{i}\right)$
@@
17:        $\theta\leftarrow\text{MergeWeight}(\theta_{\text{pre}},\mathcal{T}^{\prime\prime};\lambda^{\prime})$ // Update the merged model with the updated $\lambda^{\prime}$
@@
18:        $\Delta\leftarrow\Delta-\eta_{2}\nabla_{\Delta}\left(\sum_{i=1}^{n}l_{i}^{perturbation}\right)$ // Adversarial Trigger Recovery
19:     end if
20:     for each task $s_{i}\in\mathcal{S}$ do
21:        Sample a batch of unlabeled data $\mathcal{D}_{i}$ from $s_{i}$
22:        $l_{i}\leftarrow\mathcal{L}_{i}(f(\theta),\mathcal{D}_{i})$
23:        $l_{i}^{perturbation}\leftarrow\mathcal{L}_{i}(f(\theta),\mathcal{D}_{i}+\Delta_{i})$
24:     end for
25:     $\mathbf{x}\leftarrow\mathbf{x}-\eta_{3}\nabla_{\mathbf{x}}\left(\sum_{i=1}^{n}(l_{i}+\alpha l_{i}^{perturbation})\right)$
26:  end for
27:  Return: the mask parameterized by logits $\mathbf{x}$.
Corollary 1 (Performance-safety trade-off control).
B.4 Discussions with the robustness of backdoor attack, fine-tuning, and continual learning methods
We would like to clarify that our setting is indeed different from your mentioned robustness of backdoor attacks, model fine-tuning, and continual learning methods. The core theme of our work is related to model merging using different models (partial backdoor) rather than one model like your mentioned works.
For the traditional backdoor attack, the model provided by the adversary is the final deployment model. However, for model merging, the adversary only contributes to parts of the models, which are provided for the latter model merging, and the adversary has blind knowledge about how model merging is conducted (Zhang et al., 2024). We are the first to explore the backdoor effect (backdoor succession and backdoor transfer) during model merging and provide a defense-aware merging method to mitigate this issue. Exactly, the object of the Backdoor Detection Mask is the same as existing trigger inversion or synthesis methods (Sun & Kolter, 2023; Dunnett et al., 2024), aiming to find a backdoor trigger inserted into the model. The key difference among them exists in the optimization process and special optimization constraints.
Moreover, some trigger inversion works need to recover the backdoor through an optimization process to flip a support set of clean images into the target class (e.g.smoothinv (Sun & Kolter, 2023)) and other works (e.g. BEAGLE (Dunnett et al., 2024)) propose model backdoor forensics techniques and need a few attack samples as instructions. For our proposed DAM as shown in Table 1, the optimization of the backdoor detection mask only needs unlabeled test data. Simultaneously, as shown in Figure 4, both the Backdoor Detection Mask and the Task-Shared Mask contribute to the whole merging process and they are optimized alternately in an iterative process to develop a merged model that effectively balances performance and safety.
Moreover, model merging has its unique challenges compared with finetuning-based methods and continuous learning methods (Zhu et al., 2024; 2025). From the perspective of the problem: Fine-tuning-based methods (Zhu et al., 2023; Huang et al., 2024) directly add perturbations to the final model during fine-tuning, continual learning methods additionally consider the forgetting issues related to backdoor attacks during sequential training (Mi et al., 2023; Abbasi et al., 2024; Guo et al., 2024), but both of them only focus on the optimization of its single model during training. In contrast, model merging should additionally consider the interference from other task-specific models. As shown in Figure 4, masking the backdoor-related parameters will also influence the parameters of task interference. There exists a trade-off between performance and safety due to the conflict of these two parts of parameters, which is special for model merging. From the perspective of training Data available: As shown in Table 1, we only have the unlabeled test data for model merging, but for finetuning-based and continual learning methods, they need some labeled data at least, which means the setting of model merging is different and difficult.
B.5 Discussions about the parameter disentanglement for task vectors
To clarify our contribution more clearly, we can disentangle the parameter components for the task vectors into two parts: task-general and task-specific parts. The task-general part represents the common parts for different task-specific models. When merging the parameters of clean and backdoored task-specific models:
(i) If the backdoor is injected in the task-general region on the backdoored model, for model merging(average the model weights from clean models and backdoored models), this backdoor effect can be seen as transferring from backdoored models to clean models. That’s why the SVHN’s ASR significantly increases after merging with the clean model. It’s a unique and special phenomenon of existing model merging, which aims to utilize existing checkpoints to construct a new model without the training data for these checkpoints.
(ii) Moreover, if the backdoor is injected in the task-specific region on the backdoor model, the solution can be seen as similar to traditional backdoor defense works, because we don’t need to consider the impact of the backdoored models on other clean models.
Exactly, for defenders, we only have different checkpoints without the knowledge of whether or not and which region they are injected with the backdoor. To solve the (i) and (ii) simultaneously, our proposed DAM design two masks to identify the parameters and reset them to pre-trained weights to solve the problem, with the assumption that the pre-trained model is clean and protected.
@@ C.2 More Experiments about Backdoor Succession and Backdoor Transfer
From the results, we can observe that the evaluation of existing merging methods is not fully consistent with the results shown in (Tang et al., 2024a) when we take performance and safety into consideration simultaneously. For example, Ties-Merging achieved unexpectedly poor results, sometimes even worse than RegMean and Fisher Merging, which is the opposite of its reported outcome. This can be attributed to the fact that the success of Ties-Merging relies heavily on the gradient conflict analysis of different tasks. However, this analysis just considers the task-specific performance without dealing with safety issues. When there exists backdoored task-specific models during merging, the causes of gradient conflicts are complex and multidimensional. Notably, compared with task-wise methods, Adamerging(task-wise) and Concrete AM(task-wise), the layer-wise version of corresponding methods can consistently achieve better performance, but their ASR results are still high. This further clarifies the single task-wise performance perspective is not always appropriate, highlighting the need for more exploration of the backdoor issues during multi-task merging.
@@ C.3 More Merging Results
Moreover, Table 11 and Table 12 display the merging results while merging two backdoored models and four clean models. Different from Table 2 in the main text, we select other task-specific backdoored models during merging. These results can further verify the effectiveness and robustness of the proposed DAM.
C.4 Efficiency Studies
Table 8: Efficiency studies for the comparisons between DAM and post-defense methods.
ANP AWM SAU DAM
Merging Time 29 29 29 /
Post-Defense time 15 18 20 /
Total Time 44 47 49 35
Exactly, the reported post-defense methods (AWM and SAU) are two-stage methods (merge first and then defense), and our proposed DAM is an end-to-end merging method without post-hoc cost (consider safety during merging). We can achieve comparable or better performance compared with previous post-defense methods without additional training. We also report the training time to further clarify our contributions, where we calculate the train time (minutes) to achieve the merged model using six-task specific models considering the safety issues on a single Tesla V100 GPU with 32G memory (set the AdamW as the optimizer and the batch size as 16).
C.5 The Effect of Domain Source on Model Merging
It’s worthwhile to discuss the domain source of the models used for model merging, which has been neglected by all previous merging methods. To find out the relationship between ASR drop and domain distribution. We first carefully review six used image classification datasets into four categories from the perspective of domain source: (i) Digit images: MNIST and SVHN;(ii) Remote sensing images: RESISC45 and EUROSAT; (iii) Texture images: DTD; (iv) 3D Objects related cars: Stanford cars.
Then, we provide the merging experiments when only the task-specific model on DTD is injected with the backdoor. Other task-specific models are clean and have different domains from the task-specific model on DTD. In this way, we can conduct merging experiments when one backdoor model in a certain domain + several models (backdoored or clean) from different domains. From the results shown in Table 15, we can find that previous merging methods can naturally weaken that backdoor effect through parameter-level merging. But this doesn’t verify that our work about defense-aware merging is meaningless. There are two reasons as follows: (i) Take the results on AdaMerging (Layer-Wise) for example, though the ASR on the DTD can decrease badly after merging(98.90->31.33), the ASR on other tasks such as EUROAST(24.37-56.56) can increase unexpectedly. This means apart from the backdoor effect on the task related to the injected model, we should also focus on the ASR on the task related to the clean model used for merging. This new phenomenon during model merging has been explained as the Backdoor Transfer; (ii) Our proposed DAM further lowers the ASR compared with previous merging methods while sacrificing only about 1 in accuracy, achieving the best trade-off between performance and safety.
Moreover, in a way, introducing additional clean models for backdoored models on the same task can be seen as exploring the effect of merging models from the same domain. These experimental results can be also found in Table 5 in the paper. Specifically, during model merging, we only have different checkpoints without the knowledge of whether they are injected with a backdoor or not. Thus, in our paper, it’s reasonable to explore whether directly merging the clean models and backdoored model on the same task is enough to mitigate the backdoor as you said. For each task related to the backdoored individual finetuned model, we additionally select a clean model for this task during multi-task merging. Notable, both WAG and LoRA-as-an-Attack defend the backdoor by directly averaging the homogeneous clean and backdoored full model weights or LoRa, we implement them by averaging the weights of original task-specific models and additionally introduced clean models. From the results, we can observe that DAM consistently achieves higher ACC and lower ASR in different settings. These results can verify that the backdoor effect from task-specific models can be mitigated by the clean model from the same domain in a way, but our proposed DAM further achieves higher ACC and lower ASR in different settings.
[…]